python #Ignore line 2, it is just there to align with zybooks requirements print("5") #Create a...

python

#Ignore line 2, it is just there to align with zybooks requirements
print("5")
#Create a list with the following numbers and assign it to a variable name: 1,2,3,4,5,6
#Convert the list to a tuple and assign it to a variable name (dirrerent from the list's name)
#Try the following and press the run program button to which ones work and which ones cause errors
#If a statement causes an error, use a # to comment it out
#Add the number 7 to the beggining of the list
#Add the number 7 to the beginning of the tuple
#Append the number 8 to the end of the list
#Append the number 8 to the end of the tuple
#Remove the second element from the list using the pop method
#Remove the second element from the tuple using the pop method
#Get the sum and length of the list and compute the average of the elements in the list
#Get the sum and length of the tuple and compute the average of the elements in the tuple
#Create a tuple with the following numbers and assign it to a new variable name: 2,4,6,8,9
#Convert the tuple to a list and assign it to a new variable name
Answer #1

NOTE: For making it easy to undestand , i first uploaded the parts of the code and with outputs and then submitted the overall code at the end

"HERE IN THE GIVEN STATEMENTS"

"EXCEPT "APPENDING" AND "POPING" FOR TUPLE, ALL THE OTHER STATEMENTS GIVE OUTPUTS"

"appending and poping of tuple doesn't give outputs because "tuples" are immutable"


code for creating a list and converting it into a tuple, adding 7 and appending 8 in a list and tuple

lis=[1,2,3,4,5,6]
tup=tuple(lis)
n=7
lis=[n]+lis
print(lis)
tup=(7,)+tup
print(tup)
lis.append(8)
print(lis)
#tup.append(8)
#we cant use append in a tuple because they are immutable


proof with output:

code for removing second element and finding sum , length , average of list and tuple

lis=[1,2,3,4,5,6]
tup=tuple(lis)
lis.pop(1)
print(lis)
#tup.pop(1)
#we cant use pop in tuples because they are immutable
length=len(lis)
s=0
for l in lis:
    s=l+s
average = s / len(lis)
print(length,s,average)
length=len(tup)
s=0
for l in tup:
    s=l+s
average = s / len(tup)
print(length,s,average)

proof with output:

code for creating a new tuple and converting it into a list and assigning new variables to it:

lis=[1,2,3,4,5,6]
tup=tuple(lis)
newtup=(2,4,6,8,9)
print(newtup)
print(type(newtup))
lis2=list(newtup)
print(lis2)

proof with output:

overall code for the given question:

lis=[1,2,3,4,5,6]
tup=tuple(lis)
n=7
lis=[n]+lis
print(lis)
tup=(7,)+tup
print(tup)
lis.append(8)
print(lis)
#tup.append(8)
#we cant use append in tuple
lis.pop(1)
print(lis)
#tup.pop(1)
#we cant use pop in tuple
length=len(lis)
s=0
for l in lis:
    s=l+s
average = s / len(lis)
print(length,s,average)
length=len(tup)
s=0
for l in tup:
    s=l+s
average = s / len(tup)
print(length,s,average)
newtup=(2,4,6,8,9)
print(newtup)
print(type(newtup))
lis2=list(newtup)
print(lis2)

proof with output: